OpenStudy (shaleiah):

(3x^4 + 9x^3 – 7x + 15) + (–6x^4 – 8x^2 + 5x – 3). Help me solve (simple way)

OpenStudy (shaleiah):

@FortyTheRapper

OpenStudy (calculusxy):

Combine the like terms. Like terms share the same variable raised to the same power.

OpenStudy (shaleiah):

3x^4-6x^4

OpenStudy (calculusxy):

yes that's one.

OpenStudy (skullpatrol):

because 3x^4-6x^4 = ( 3 - 6 ) x^4

OpenStudy (shaleiah):

9x^3+(-7x+15)

OpenStudy (calculusxy):

these are not like terms

OpenStudy (skullpatrol):

recall ba + ca = ( b + c ) a

OpenStudy (skullpatrol):